# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

### v2

@@ 1 Introduction
Another challenge limiting the impact of public training data is that most datasets are computed using relatively low-level QM methods, with DFT being the most common method of choice. [22] Although modern functionals can achieve reasonable accuracy given a large enough basis set, most QM datasets are computed with traditional functionals and moderate basis sets. Unfortunately, highly accurate QM methods (e.g., coupled-cluster, CC, [24] with a complete basis set extrapolation) [25] are orders of magnitude more computationally expensive. Consequently, the few datasets computed with high-level QM methods are much smaller in terms of both number of calculations and size of systems (i.e., number of atoms). [17]
@@
The abundance of incompatible low-accuracy datasets and the scarcity of high-accuracy forces requires more sophisticated training techniques that can simultaneously learn from multiple datasets. Figure 1 compares four such multi-dataset training paradigms, namely, transfer learning (TL), [32, 37, 39, 40] $\Delta$-learning ($\Delta$L), [31, 33, 41, 42, 43] meta-learning ($\mu$L), [44] and multi-fidelity learning (MFL). [45, 46] All four methods present unique advantages and disadvantages when training to multiple datasets with different levels of theory. Both TL and $\Delta$L are powerful methods in the limit of only two datasets at different levels of theory, whereas $\mu$L and MFL are theoretically amenable to any number of levels of theory. Both TL and $\Delta$L consist of first training an MLIP on the large, low-accuracy dataset, followed by training to the smaller, high-accuracy dataset. By contrast, $\mu$L and MFL train to all datasets simultaneously, although $\mu$L also requires a subsequent fine-tuning step. Thus, $\Delta$L, and $\mu$L are two-step methods that result in separate MLIPs for each level of theory. By contrast, MFL is a single-step method which trains a single MLIP that predicts multiple energies, with one output energy for each level of accuracy (i.e., level of fidelity). MFL is also often referred to as multi-level learning, [47] multi-head learning, and multi-task learning [48, 49] (see Appendix for discussion of terminology).
Figure 1: Comparison of different methods capable of training to multiple datasets at different levels of theory. Although meta-learning and multi-fidelity learning can utilize many datasets, for visual simplicity, only three levels of theory are depicted. Low-level data are represented with density functional theory (DFT). Intermediate-level data are represented with Møller-Plesset second-order perturbation theory (MP2). High-level data are represented with coupled-cluster theory (CC). While some variation exists in the literature as to the implementation of each method, the depictions in this figure are consistent with the TL and $\Delta$L work of Smith et al., [37] the $\mu$L work of Allen et al., [44] and the MFL work of Jacobson et al. [45]
Previous studies demonstrate that TL, [37] $\Delta$L, [37] and $\mu$[44] achieve significant improvement when training to two datasets that overlap perfectly in configuration space, namely, the ANI-1x and ANI-1ccx datasets. Specifically, all three multi-dataset methods report approximately 20% higher accuracy compared to training solely to the ANI-1ccx CC-level energies. By contrast, previous MFL work did not see significant improvement when simultaneously training to ANI-1x and ANI-1ccx. [45] However, this study did not include the DFT-level atomic forces when performing MFL, only the DFT-level energies and CC-level energies were used for training. The justification for training only to energies was to avoid biasing the MLIP towards the ANI-1x DFT-level forces, since ANI-1ccx does not contain CC-level force data. Consequently, MFL with just DFT-level energies and CC-level energies achieved comparable accuracy to training solely to CC-level energies. Additional studies also tested the limits of MFL by training to several different levels of theory simultaneously. [45, 47, 50, 51] Learning from multiple datasets covering different regions of chemical space offers clear benefits for improved transferability, but does not necessarily lead to improved accuracy.
The primary goal of this work is to demonstrate that MFL can result in an MLIP with extremely accurate high-level energies and forces, even when high-level forces are not available for training. Our hypothesis is that high-level forces can be learned implicitly with MFL by simultaneously training to low-level forces and high-level energies. To validate our hypothesis, we apply MFL to train an MLIP to the DFT-level energies and forces and CC-level energies for the $\approx$460k structures common to both the ANI-1x and ANI-1ccx datasets. These two levels of theory were chosen to enable comparison with previous studies that trained to the same datasets but utilized TL, [37] $\Delta$L, [37] $\mu$L, [44] and MFL without DFT-level forces. [45] For further validation, we repeat this MFL process but with two other intermediate-level QM methods in place of the CC-level energy data. For completeness, we briefly investigate any potential benefit of training with force data from more than two levels of theory.
@@ 2.2.1 Training
The reduced ANI-1x dataset contains energies and forces computed with several QM methods. To simplify our analysis, we focus on only four levels of theory: DFT/DZ, DFT/TZ, MP2/TZ, and CCSD(T)*/CBS. The two lowest fidelity levels (DFT/DZ and DFT/TZ) are based on the hybrid $\omega$b97x functional [62] with either the 6-31G* double zeta (DZ) basis set [63] or the def2-TZVPP triple zeta (TZ) basis set. [64] For brevity, we refer to these two DFT-based fidelity levels simply as DZ and TZ, respectively. The second highest fidelity level (MP2/TZ) corresponds to Møller-Plesset second-order perturbation theory (MP2) [65] with the cc-pVTZ triple zeta basis set. [66, 67] The highest fidelity level (CCSD(T)*/CBS) utilizes an approximate coupled-cluster (CC) single and double excitation with perturbative triple excitation [68] with a complete basis set (CBS) extrapolation implementation [69, 70] (see Smith et al. [17] for details). For brevity, we refer to the two highest fidelity levels as MP2 and CC, respectively. While energies are available for nearly all $\approx$460k structures across all four fidelity levels, forces are only available for the DFT/DZ and DFT/TZ methods.
@@ 2.2.2 Testing
An additional out-of-sample test dataset contains torsion scan energy differences $(\Delta E_{\rm tors})$ for small drug-like molecules. This test dataset contains the same geometries as Sellers et al. [74] that were previously optimized with MP2 and the 6-311+G** triple zeta basis set. [75] However, the energies in our torsion scan test dataset were previously recomputed using the same CCSD(T)*/CBS approach as the ANI-1ccx dataset. We also test our models against the CCSD(T)/CBS energies reported by Sellers et al. (see Supporting Information), to allow for a fair comparison with other studies. [37, 45, 58, 76] To further test the robustness of our models, we also perform relaxed torsion scans, wherein the positions of all atoms are optimized with a given model subject to the constraint of a fixed dihedral angle (see Supporting Information).
@@ 2.3 Models
A comprehensive set of training tests is performed with each level of theory for both single-fidelity (SF) and multi-fidelity (MF) machine learning interatomic potentials (MLIPs). Table 1 gives a summary of all MLIPs trained in this work, as well as the shorthand notation adopted to distinguish between models. The SF-MLIPs allow for a baseline comparison with the MF-MLIPs as well as a benchmark comparison with the published ANI MLIPs trained to the same level of theory. An ensemble consisting of eight MLIPs is trained for each model. See Supporting Information for model and training details.
Table 1: All MLIPs trained in this work. SF = single-fidelity. MF = multi-fidelity. DZ = DFT/DZ. TZ = DFT/TZ. MP2 = MP2/TZ. CC = CCSD(T)*/CBS. E = energies. F = forces. ✓signifies MLIP was trained with data from this level of theory (DFT/DZ, DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) and property (E or F).
@@ 3.1 Test case one: DFT/DZ and DFT/TZ
Table 2 reports the energy and force RMSEs for an SF-MLIP trained just to DFT/TZ energies (SF-TZ/E), an SF-MLIP trained to DFT/TZ energies and forces (SF-TZ/EF), an MF-MLIP trained to DFT/DZ energies and DFT/TZ energies (MF-DZ/E-TZ/E), an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies (MF-DZ/EF-TZ/E), and an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies and forces (MF-DZ/EF-TZ/EF). See Supporting Information for additional validation results relevant to the DFT/DZ and DFT/TZ test case.
@@ 3.2 Test case two: DFT/DZ and MP2/TZ
We now investigate MFL with DFT/DZ and MP2/TZ data. Although the training dataset does not contain any MP2/TZ force data, this test case allows for validation against the MP2/TZ forces in the GDB10to13 dataset. In the first test case, the two levels of theory (DFT/DZ and DFT/TZ) are relatively similar and computationally inexpensive. In this test case, however, the MP2/TZ forces are quite different from the DFT/DZ forces (RMSE of 5.9 kcal/mol/Å when comparing DFT/DZ and MP2/TZ forces in GDB10to13). The more expensive MP2/TZ forces are also assumed to be more accurate than DFT/TZ, considering MP2 is a higher level of theory and both methods use a triple zeta basis set. Therefore, this test case presents a more interesting and practical challenge for MFL than test case one.
Table 3 reports the energy and force RMSEs for a single-fidelity MLIP trained just to MP2/TZ energies (SF-MP2/E) and a multi-fidelity MLIP trained to DFT/DZ energies and forces and MP2/TZ energies (MF-DZ/EF-MP2/E). See Supporting Information for additional validation results relevant to the DFT/DZ and MP2/TZ test case.
@@ 3.3 Test case three: DFT/DZ and CCSD(T)*/CBS
Having verified that MFL quantifiably improves the high-level energies and forces for the first two test cases, we now aim at the most challenging and important test case, namely, MFL with DFT/DZ and gold-standard CCSD(T)*/CBS data. Because CC is prohibitively expensive, no CC-level force data are available for direct validation in either the ANI-1ccx training dataset or the GDB10to13 test dataset. However, Supporting Information provides indirect evidence that MFL with low-level forces improves CC-level forces.
@@
Table 4 reports the energy and $\Delta E_{\rm conf}$ RMSEs for a single-fidelity MLIP trained to CCSD(T)*/CBS energies (SF-CC/E), a multi-fidelity MLIP trained to DFT/DZ energies and CCSD(T)*/CBS energies (MF-DZ/E-CC/E), and a multi-fidelity MLIP trained to DFT/DZ energies and forces and CCSD(T)*/CBS energies (MF-DZ/EF-CC/E). See Supporting Information for additional validation results relevant to the DFT/DZ and CCSD(T)*/CBS test case.
@@
Similar to the results for the previous two test cases, the results in Table 4 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (CCSD(T)*/CBS) energies. Specifically, the MF-DZ/EF-CC/E errors are considerably lower than the SF-CC/E errors on CCSD(T)*/CBS energies (1.52 kcal/mol vs 2.78 kcal/mol), CCSD(T)*/CBS conformer energy differences (1.60 kcal/mol vs 2.65 kcal/mol), and CCSD(T)*/CBS torsion scan energy differences (0.40 kcal/mol vs 0.71 kcal/mol), respectively. The MF-DZ/EF-CC/E energy, $\Delta E_{\rm conf}$, and $\Delta E_{\rm tors}$ RMSEs (1.52 kcal/mol, 1.60 kcal/mol, and 0.40 kcal/mol) are also substantially lower than the RMSEs for the ANI-1ccx MLIP (2.57 kcal/mol, 2.62 kcal/mol, and 0.77 kcal/mol), respectively. See Supporting Information for additional comparisons between HIP-NN and ANI models. By contrast, MFL without forces (i.e., only energies for both levels of theory) is essentially as accurate as SF training to just the high-level energies, consistent with previous work. [45] This result demonstrates that including low-level forces is essential for MFL to achieve significant improvement in accuracy.
@@ 4 Discussion
These three test cases clearly demonstrate that MFL with low-level forces and high-level energies outperforms SF training with only high-level energies. Moreover, MFL with low-level forces is of similar accuracy as SF training directly to high-level forces, in cases where those are available for comparison. To help quantify the improvement MFL provides in all three test cases, Figure 3 presents the percent decrease in RMSE for energies, forces, and conformer energy differences between each MF-MLIP trained with low-level (DFT/DZ) forces and the corresponding SF-MLIP trained with only high-level (either DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) energies. See Supporting Information for additional analysis of percent improvement.
@@
The percent improvement in energy and conformer energy differences with MFL is between 40% and 45% for all three test cases. The percent improvement in forces with MFL is 70% and 60% for MF-DZ/EF-TZ/E tested on the ANI-1x DFT/TZ forces and MF-DZ/EF-MP2/E tested on the GDB10to13 MP2/TZ forces, respectively. By contrast, the percent improvement with transfer learning (TL) reported for the ANI-1ccx MLIP is considerably smaller (between 15-20% for energies and conformer energies) than with any of the MFL test cases. For a truly fair comparison between TL and MFL, however, future work could investigate the percent improvement for TL that is achieved when pre-training with the same HIP-NN architecture and the same reduced ANI-1x dataset with both DFT/DZ energies and DFT/DZ forces.
@@
One unexpected result is that MFL leads to a measurable decrease in accuracy on the low-level forces themselves (see Supporting Information). Specifically, the RMSEs on the DFT/DZ forces for MF-DZ/EF-TZ/E (3.95 kcal/mol/Å), MF-DZ/EF-MP2/E (3.91 kcal/mol/Å), and MF-DZ/EF-CC/E (3.76 kcal/mol/Å) are all markedly higher than the force RMSE for SF-DZ/EF (3.17 kcal/mol/Å). This increase in error on the low-level (DFT/DZ) forces suggests that the current MFL approach requires a compromise in accuracy on the low-level forces to achieve improved accuracy on the high-level forces. Future work could attempt to elucidate whether this is a fundamental limitation of MFL, a limitation of our implementation, or a limitation in the model architecture. Some possible considerations to help reduce the low-level force errors are to use a larger HIP-NN architecture (i.e., more interactions layers, more atom layers, more features), and/or modify the number of shared parameters, and/or include non-linear layers for the fidelity-specific output nodes. However, we emphasize that the primary goal of this work is to improve the high-level energy and force predictions. Indeed, if only accurate prediction of the low-level energy and forces is the desired outcome, SF learning would suffice.
@@
For completeness, we also investigate the potential benefit of applying MFL with forces from more than one level of theory. In this case, the energies and forces for both DFT/DZ and DFT/TZ are included in training along with the CCSD(T)*/CBS energies. Notably, the CC-level errors are actually considerably worse for MFL with three datasets than MFL with just two datasets (see Supporting Information). With twice as much DFT-level data, the worse performance with three datasets is likely attributed to a data imbalance. An alternative explanation is that the DFT/DZ and DFT/TZ forces are providing contradictory information that MFL simply cannot resolve. Another possibility is that the MLIP has reached its capacity to match the different datasets with the chosen architecture and hyperparameters. However, with the current MFL approach, low-level forces from just a single level of theory are sufficient to achieve the desired accuracy on the high-level forces.

### v3

@@ 1 Introduction
Another challenge limiting the impact of public training data is that most datasets are computed using relatively low-level QM methods, with DFT being the most common method of choice. [22] Although modern functionals can achieve reasonable accuracy given a large enough basis set, most QM datasets are computed with traditional functionals and moderate basis sets. Unfortunately, highly accurate QM methods (e.g., coupled-cluster, CC, [24] with a complete basis set, CBS, extrapolation) [25] are orders of magnitude more computationally expensive. Consequently, the few datasets computed with high-level QM methods are much smaller in terms of both number of calculations and size of systems (i.e., number of atoms). [17]
@@
Force data are especially valuable for MLIP development. Because forces are vector quantities associated with each atom, a single QM calculation returns one system energy but 3NN force components, where $N$ is the number of atoms. Forces are also a local property, whereas system energy is a global property. Thus, force data facilitate more direct learning of the relationship between the local atomic environment and the potential energy surface. Furthermore, as forces are the negative gradient of the energy, training to forces helps ensure a “smooth” potential. For these reasons, inclusion of force data [34] (and, by extension, Hessian data) [35, 36] is extremely advantageous for training speed, MLIP accuracy, and MLIP stability in MD simulations. For example, due to both the lack of force data in ANI-1ccx and the fewer number of molecular structures in ANI-1ccx compared to ANI-1x, previous MLIPs trained solely to the ANI-1ccx dataset were more limited in accuracy despite being trained to more expensive CC-level energy data. [37, 38, 39]
@@
Figure 1: Comparison of different methods capable of training to multiple datasets at different levels of theory. Although meta-learning and multi-fidelity learning can utilize many datasets, for visual simplicity, only three levels of theory are depicted. Low-level data are represented with density functional theory (DFT). Intermediate-level data are represented with Møller-Plesset second-order perturbation theory (MP2). High-level data are represented with coupled-cluster theory ($C$). While some variation exists in the literature as to the implementation of each method, the depictions in this figure are consistent with the transfer learning and $\Delta$-learning work of Smith et al., [37] the meta-learning work of Allen et al., [44] and the multi-fidelity learning work of Jacobson et al. [45]
[…]
@@ 2.2.1 Training
[…]
@@ 2.2.2 Testing
An additional out-of-sample test dataset contains torsion scan energy differences $(\Delta E_{\rm tors})$ for small drug-like molecules. This test dataset contains the same geometries as Sellers et al. [74] that were previously optimized with MP2 and the split-valence TZ with diffuse and polarization functions (6-311+G**) basis set. [75] However, the energies in our torsion scan test dataset were previously recomputed using the same CCSD(T)*/CBS approach as the ANI-1ccx dataset. We also test our models against the CCSD(T)/CBS energies reported by Sellers et al. (see Supplementary Material), to allow for a fair comparison with other studies. [37, 45, 59, 76] To further test the robustness of our models, we also perform relaxed torsion scans, wherein the positions of all atoms are optimized with a given model subject to the constraint of a fixed dihedral angle (see Supplementary Material).
@@ 2.3 Models
A comprehensive set of training tests is performed with each level of theory for both single-fidelity (SF) and multi-fidelity (MF) machine learning interatomic potentials (MLIPs). Table 1 gives a summary of all MLIPs trained in this work, as well as the shorthand notation adopted to distinguish between models. The SF-MLIPs allow for a baseline comparison with the MF-MLIPs as well as a benchmark comparison with the published ANI MLIPs trained to the same level of theory. An ensemble consisting of eight MLIPs is trained for each model. See Supplementary Material for model and training details.
Table 1: All MLIPs trained in this work. SF = single-fidelity. MF = multi-fidelity. DZ = DFT/DZ. TZ = DFT/TZ. MP2 = MP2/TZ. $C$ = CCSD(T)*/CBS. E = energies. F = forces. ✓signifies MLIP was trained with data from this level of theory (DFT/DZ, DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) and property (E or F).
@@ 3.1 Test case one: DFT/DZ and DFT/TZ
Table 2 reports the energy and force RMSEs for an SF-MLIP trained just to DFT/TZ energies (SF-TZ/E), an SF-MLIP trained to DFT/TZ energies and forces (SF-TZ/EF), an MF-MLIP trained to DFT/DZ energies and DFT/TZ energies (MF-DZ/E-TZ/E), an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies (MF-DZ/EF-TZ/E), and an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies and forces (MF-DZ/EF-TZ/EF). See Supplementary Material for additional validation results relevant to the DFT/DZ and DFT/TZ test case.
@@ 3.2 Test case two: DFT/DZ and MP2/TZ
We now investigate MFL with DFT/DZ and MP2/TZ data. Although the training dataset does not contain any MP2/TZ force data, this test case allows for validation against the MP2/TZ forces in the GDB10to13 dataset. In the first test case, the two levels of theory (DFT/DZ and DFT/TZ) are relatively similar and computationally inexpensive. In this test case, however, the MP2/TZ forces are quite different from the DFT/DZ forces (RMSE of 5.9 kcal/mol/Å when comparing DFT/DZ and MP2/TZ forces in GDB10to13). The more expensive MP2/TZ forces are also assumed to be more accurate than DFT/TZ, considering MP2 is a higher level of theory and both methods use similar triple zeta basis sets. Therefore, this test case presents a more interesting and practical challenge for MFL than test case one.
Table 3 reports the energy and force RMSEs for a single-fidelity MLIP trained just to MP2/TZ energies (SF-MP2/E) and a multi-fidelity MLIP trained to DFT/DZ energies and forces and MP2/TZ energies (MF-DZ/EF-MP2/E). See Supplementary Material for additional validation results relevant to the DFT/DZ and MP2/TZ test case.
@@ 3.3 Test case three: DFT/DZ and CCSD(T)*/CBS
Having verified that MFL quantifiably improves the high-level energies and forces for the first two test cases, we now aim at the most challenging and important test case, namely, MFL with DFT/DZ and gold-standard CCSD(T)*/CBS data. Because CC is prohibitively expensive, no CC-level force data are available for direct validation in either the ANI-1ccx training dataset or the GDB10to13 test dataset. However, Supplementary Material provides indirect evidence that MFL with low-level forces improves CC-level forces.
@@
Table 4 reports the energy and $\Delta E_{\rm conf}$ RMSEs for a single-fidelity MLIP trained to CCSD(T)*/CBS energies (SF-CC/E), a multi-fidelity MLIP trained to DFT/DZ energies and CCSD(T)*/CBS energies (MF-DZ/E-CC/E), and a multi-fidelity MLIP trained to DFT/DZ energies and forces and CCSD(T)*/CBS energies (MF-DZ/EF-CC/E). See Supplementary Material for additional validation results relevant to the DFT/DZ and CCSD(T)*/CBS test case.
@@
Similar to the results for the previous two test cases, the results in Table 4 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (CCSD(T)*/CBS) energies. Specifically, the MF-DZ/EF-CC/E errors are considerably lower than the SF-CC/E errors on CCSD(T)*/CBS energies (1.52 kcal/mol vs 2.78 kcal/mol), CCSD(T)*/CBS conformer energy differences (1.60 kcal/mol vs 2.65 kcal/mol), and CCSD(T)*/CBS torsion scan energy differences (0.40 kcal/mol vs 0.71 kcal/mol), respectively. The MF-DZ/EF-CC/E energy, $\Delta E_{\rm conf}$, and $\Delta E_{\rm tors}$ RMSEs (1.52 kcal/mol, 1.60 kcal/mol, and 0.40 kcal/mol) are also substantially lower than the RMSEs for the ANI-1ccx MLIP (2.57 kcal/mol, 2.62 kcal/mol, and 0.77 kcal/mol), respectively. See Supplementary Material for additional comparisons between HIP-NN and ANI models. By contrast, MFL without forces (i.e., only energies for both levels of theory) is essentially as accurate as SF training to just the high-level energies, consistent with previous work. [45] This result demonstrates that including low-level forces is essential for MFL to achieve significant improvement in accuracy.
@@ 4 Discussion
These three test cases clearly demonstrate that MFL with low-level forces and high-level energies outperforms SF training with only high-level energies. Moreover, MFL with low-level forces is of similar accuracy as SF training directly to high-level forces, in cases where those are available for comparison. To help quantify the improvement MFL provides in all three test cases, Figure 3 presents the percent decrease in RMSE for energies, forces, and conformer energy differences between each MF-MLIP trained with low-level (DFT/DZ) forces and the corresponding SF-MLIP trained with only high-level (either DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) energies. See Supplementary Material for additional analysis of percent improvement.
@@
The percent improvement in energy and conformer energy differences with MFL is between 40% and 45% for all three test cases. The percent improvement in forces with MFL is 70% for MF-DZ/EF-TZ/E tested on the ANI-1x DFT/TZ forces and 60% for MF-DZ/EF-MP2/E tested on the GDB10to13 MP2/TZ forces. For comparison, the percent improvement with TL is more than a factor of two smaller than with any of the MFL test cases. Specifically, based on the reported errors for the ANI-1ccx and ANI-1ccx-R MLIPs [37], which were trained with and without TL, respectively, the percent improvement with TL is between 15-20% for energies and conformer energies. For a truly fair comparison between TL and MFL, however, future work could investigate the percent improvement for TL that is achieved when pre-training with the same HIP-NN architecture and the same reduced ANI-1x dataset with both DFT/DZ energies and DFT/DZ forces.
@@
One unexpected result is that MFL leads to a measurable decrease in accuracy on the low-level forces themselves (see Supplementary Material). Specifically, the RMSEs on the DFT/DZ forces for MF-DZ/EF-TZ/E (3.95 kcal/mol/Å), MF-DZ/EF-MP2/E (3.91 kcal/mol/Å), and MF-DZ/EF-CC/E (3.76 kcal/mol/Å) are all markedly higher than the force RMSE for SF-DZ/EF (3.17 kcal/mol/Å). This increase in error on the low-level (DFT/DZ) forces suggests that the current MFL approach requires a compromise in accuracy on the low-level forces to achieve improved accuracy on the high-level forces. Future work could attempt to elucidate whether this is a fundamental limitation of MFL, a limitation of our implementation, or a limitation in the model architecture. Some possible considerations to help reduce the low-level force errors are to use a larger HIP-NN architecture (i.e., more interactions layers, more atom layers, more features), and/or modify the number of shared parameters, and/or include non-linear layers for the fidelity-specific output nodes. However, we emphasize that the primary goal of this work is to improve the high-level energy and force predictions. Indeed, if only accurate prediction of the low-level energy and forces is the desired outcome, SF learning would suffice.
@@
For completeness, we also investigate the potential benefit of applying MFL with forces from more than one level of theory. In this case, the energies and forces for both DFT/DZ and DFT/TZ are included in training along with the CCSD(T)*/CBS energies. Notably, the CC-level errors are actually considerably worse for MFL with three datasets than MFL with just two datasets (see Supplementary Material). With twice as much DFT-level data, the worse performance with three datasets is likely attributed to a data imbalance. An alternative explanation is that the DFT/DZ and DFT/TZ forces are providing contradictory information that MFL simply cannot resolve. Another possibility is that the MLIP has reached its capacity to match the different datasets with the chosen architecture and hyperparameters. However, with the current MFL approach, low-level forces from just a single level of theory are sufficient to achieve the desired accuracy on the high-level forces.